OpenStudy (anonymous):

the area of a rectangular cloth is (6x^2-19x-85) cm^2 the length is (2x+5) cm. find the width

OpenStudy (anonymous):

\[Area = Length \times width\] \[6x^2-19x-85 = 2x+5 \times width\]